# Instruction-tuned Language Models are Better Knowledge Learners

### v1

@@ Abstract
Zhengbao Jiang††thanks: Majority of the work done during an internship at Meta. Zhiqing Sun Weijia Shi Pedro Rodriguez Chunting Zhou Graham Neubig Xi Victoria Lin Wen-tau Yih Srinivasan Iyer FAIR at Meta Carnegie Mellon University University of Washington {zhengbaj,gneubig}@cs.cmu.edu {victorialin,scottyih,sviyer}@meta.com
@@ 2.2 Wiki2023 Question-answer Pairs
To collect QA pairs for either instruction-tuning or performance evaluation, we employ publicly available LLMs to generate diverse questions and their respective answers given the article as context, following the Prompt § 2.2. On average, 4.93 questions are generated for each article. Fig. 2 and Fig. 3 show the detailed statistics and example QA pairs about “Oppenheimer”, respectively.
@@ 4.2 Analyzing the Training Dynamics: Perplexity and Generalization
How does lower perplexity of documents lead to generalization to answering related questions? We vary the number of epochs (4(a)) and learning rate (4(b)) for continued pre-training on documents and monitor three metrics to study the training dynamics.777Since we always decay the learning rate to 10% of its initial value, training for more epochs is not the same as continuing training from a checkpoint obtained after fewer epochs.
@@
As shown in 4(a), QA accuracy consistently improves as perplexity approaches one, indicating that factual knowledge learning necessitates exhaustive loss minimization over all tokens. This contrasts with learning general skills, where overly optimizing leads to overfitting.
As shown in 4(a) and 4(b), among all cases where LLMs have minimized perplexity on documents, cases trained with more epochs or larger learning rates typically exhibit superior QA performance. We hypothesize that more aggressive training leads to less overfitting to deceptive patterns in documents and better generalization when responding to questions.
@@ 5.2 Pre-instruction-tuning++
We first study how the performance varies with different numbers of epochs. As shown in Tab. 2, training for 1 epoch is insufficient, and the performance of 3, 5, or 10 epochs is similar. We fix the number of epochs to 3 and arrange the order of QA pairs and corresponding documents as shown in Fig. 6. The interleaved arrangement cycles through all the data 3 times, ensuring that in each epoch, questions either precede or follow their associated documents. On the other hand, the grouped arrangement clusters each example’s 3 appearances together, guaranteeing that the repeated questions are positioned either before or after their respective repeated documents. As shown in Tab. 2, positioning QA pairs before corresponding documents achieves better performance in both grouped and interleaved arrangements, indicating that during PIT, the learning mechanism prioritizes understanding how to access knowledge before learning to absorb information from the more complex and information-dense documents.
@@ 6.3 Analyzing the Training Dynamics of LMs
Many works study the training dynamics of LMs from different perspectives. Carlini et al. (2022) quantifies memorization across model sizes and the frequency of data duplication. Tirumala et al. (2022) finds that larger LMs memorize training data faster with less overfitting. Xia et al. (2023) show that perplexity is more predictive of model behaviors than other factors. Dery et al. (2022) studies end-task aware pre-training using classification tasks and RoBERTa models. Our work differs in that we specifically focus on the capacity of recalling and generalizing information from a seen document to answer questions.
@@ 7 Conclusion
We study the best way of continued training on new documents with the goal of later eliciting factual knowledge. We propose pre-instruction-tuning that learns how knowledge is accessed via QA pairs prior to encoding knowledge from documents. Extensive experiments demonstrate the superiority of pre-instruction-tuning versus standard instruction-tuning. Future directions include scaling this method up to a broader range of documents and instructions for more robust generalization.
Limitations

### v2

@@ Abstract
Zhengbao Jiang2††thanks: Majority of the work done during an internship at Meta. Zhiqing Sun2 Weijia Shi1,3 Pedro Rodriguez1 Chunting Zhou1 Graham Neubig2 Xi Victoria Lin1 Wen-tau Yih1 Srinivasan Iyer1 1FAIR at Meta 2Carnegie Mellon University 3University of Washington {zhengbaj,gneubig}@cs.cmu.edu {victorialin,scottyih,sviyer}@meta.com
@@ 2.2 Wiki2023 Question-answer Pairs
To collect QA pairs for either instruction-tuning or performance evaluation, we employ publicly available LLMs to generate diverse questions and their respective answers given the article as context, following the Prompt Appendix A in Appendix A. On average, 4.93 questions are generated for each article. Fig. 2 and Fig. 3 show the detailed statistics and example QA pairs about “Oppenheimer”, respectively.
@@ 3.2 Hyperparameters
When pre-training on documents, we use a batch size of 256 documents and an initial learning rate of 3e-5. During instruction-tuning on QA pairs, we use the same batch size of 256 QA pairs but opt for a reduced initial learning rate of 5e-6 because the number of tokens in a single batch is lower. Details can be found in Appendix B.
@@ 3.3 Evaluation Metrics
Since most answers are relatively short, we use exact match (EM) as our primary metric Kwiatkowski et al. (2019). To assess longer responses and accommodate minor lexical differences, we also report answer recall and ROUGE-L. Details can be found in Appendix C.
@@ 4.2 Analyzing the Training Dynamics: Perplexity and Generalization
How does lower perplexity of documents lead to generalization to answering related questions? We vary the number of epochs (5(a)) and learning rate (5(b)) for continued pre-training on documents and monitor three metrics to study the training dynamics.777Since we always decay the learning rate to 10% of its initial value, training for more epochs is not the same as continuing training from a checkpoint obtained after fewer epochs.
@@
As shown in 5(a), QA accuracy consistently improves as perplexity approaches one, indicating that factual knowledge learning necessitates exhaustive loss minimization over all tokens. This contrasts with learning general skills, where overly optimizing leads to overfitting.
As shown in 5(a) and 5(b), among all cases where LLMs have minimized perplexity on documents, for reasonably small learning rates (5e-5 is too large and leads to overfitting), cases trained with more epochs or larger learning rates typically exhibit superior QA performance. We hypothesize that more aggressive training leads to less overfitting to deceptive patterns in documents and better generalization when responding to questions.
@@ 5.2 Pre-instruction-tuning++
We first study how the performance varies with different numbers of epochs. As shown in Tab. 2, training for 1 epoch is insufficient, and the performance of 3, 5, or 10 epochs is similar. We fix the number of epochs to 3 and arrange the order of QA pairs and corresponding documents as shown in Fig. 6 in Appendix D. The interleaved arrangement cycles through all the data 3 times, ensuring that in each epoch, questions either precede or follow their associated documents. On the other hand, the grouped arrangement clusters each example’s 3 appearances together, guaranteeing that the repeated questions are positioned either before or after their respective repeated documents. As shown in Tab. 2, positioning QA pairs before corresponding documents achieves better performance in both grouped and interleaved arrangements, indicating that during PIT, the learning mechanism prioritizes understanding how to access knowledge before learning to absorb information from the more complex and information-dense documents.
@@ 6.3 Analyzing the Training Dynamics of LMs
Many works study the training dynamics of LMs from different perspectives. Carlini et al. (2022) quantifies memorization across model sizes and the frequency of data duplication. Tirumala et al. (2022) finds that larger LMs memorize training data faster with less overfitting. Xia et al. (2023) shows that perplexity is more predictive of model behaviors than other factors. Dery et al. (2022) studies end-task aware pre-training using classification tasks and RoBERTa models. Jia et al. (2022) adds a pre-training objective to encourage the vector for each phrase to have high similarity with the vectors for all questions it answers. Our work differs in that we specifically focus on the capacity of recalling and generalizing information from a seen document to answer questions.
@@ 7 Conclusion
We study the best way of continued training on new documents with the goal of later eliciting factual knowledge and propose pre-instruction-tuning that learns how knowledge is accessed via QA pairs prior to encoding knowledge from documents. Extensive experiments and ablation studies demonstrate the superiority of pre-instruction-tuning versus standard instruction-tuning. Future directions include scaling this method up to a broader range of documents and instructions for more robust generalization.
8 Limitations
@@ Acknowledgements
Appendix A Wiki2023 Dataset
Appendix B Hyperparameters
Appendix C Evaluation Metrics
Appendix D Details of Ablation Studies
We arrange the order of QA pairs and corresponding documents as shown in Fig. 6 to study the learning mechanism of pre-instruction-tuning.